OpenStudy (rachie19):

Assuming an efficiency of 31.50 % calculate the actual yield of magnesium nitrate formed from 142.4 of magnesium and excess copper (II) nitrate Mg + Cu(NO3)2 = Mg(NO3)2 +Cu

OpenStudy (rachie19):

@Rushwr

OpenStudy (rushwr):

Do u know how to get this ?

OpenStudy (rushwr):

Efficiency = actual yield/ theoretical yield *100%

OpenStudy (rushwr):

First we will have to find the theoretical yield first !

OpenStudy (rachie19):

yh i got 5.841217606g for my actual yiel

OpenStudy (rushwr):

oh let me check

OpenStudy (rachie19):

but im stuck

OpenStudy (rushwr):

where are u sucked?

OpenStudy (rachie19):

im not sure what to do next, im trying to find the theoretical yield

OpenStudy (rachie19):

i mean actual

OpenStudy (rachie19):

the one i did was the theoretical i wrote actual

OpenStudy (rushwr):

oh still wrong ! Wait. I'll tell u .

OpenStudy (rachie19):

ok thanks

OpenStudy (rushwr):

First u have to find a relationship between Mg and Mg(NO3)2 ! Cuz the information we know is about Mg right?

OpenStudy (photon336):

\(\color{blue}{\text{Originally Posted by}}\) @Rushwr First we will have to find the theoretical yield first ! \(\color{blue}{\text{End of Quote}}\) \[\frac{ actual }{ theoretical}*100 = percent, yield \]

OpenStudy (rachie19):

to get the theoretical yield. i divided the grams which was given.... 142.4/ 24.31 (1mol (NO3)2/ 1mol of Mg

OpenStudy (rachie19):

i then multiplied by 148.31 g/mol

OpenStudy (rushwr):

As we can see in the equation the stoichiometric coefficients are 1:1 in Mg:Mg(NO3)3 That means 1 mole of Mg will form Mg(NO3)2 So the next we are doing is finding the moles of Mg present. we know moles = mass divided by molar mass \[n _{Mg} = \frac{ 142.4g }{ 24gmol ^{-1} }\] Now we are gonna write the same equation for Mg(NO3)2 But here we don't know the mass, so we take the unknown mass as "x" \[n _{Mg(NO3)_{2}} = \frac{ Xg }{ 148.3gmol ^{-1} }\] As I told u earlier moles of Mg = moles of Mg(NO3)2 So we can equalize those 2 equation and find the theoretical mass of Mg(NO3)2 formed. \[\frac{ 142.4 g}{ 24gmol ^{-1} } = \frac{ X g}{ 148.3gmol ^{-1} }\] \[X= 880g \] Now this is the theoretical mass, but we have to find the actual yield . \[efficiency = \frac{ actual yield }{ Theoretical yield } * 100%\] \[actual yield = \frac{ Efficiency* theoretical yield }{ 100 } \] \[Actual yield = \frac{ 31.5 * 880 }{ 100} = 277.2g\] This is the answer I got. U better check my calculations. Others seems okai

OpenStudy (rushwr):

I hope u got it @rachie19 :)

OpenStudy (rachie19):

ok i understand what i did wrong. Thank you so much :)

OpenStudy (rushwr):

Ur welcome !!! And no problem !